OpenStudy (anonymous):

for all a,b,c are integers, there exists an integer solution to ax^2 +by^2=c, whenever gcd(a,b)|c

OpenStudy (anonymous):

can u prove this

OpenStudy (anonymous):

or disprove it

OpenStudy (anonymous):

no i cannot, at least it is not obvious

myininaya (myininaya):

i think i might be getting somewhere

OpenStudy (anonymous):

ok cuz i'm not

myininaya (myininaya):

ok let me type what i have

myininaya (myininaya):

Ok there is ordered pair (m,n) where m and n are integers such that \[am^2+bn^2=c.............(1)\] If gcd(a,b)=d, then there are integers j,k such that aj+bk=d. If gcd(a,b)|c, then there is integer l such that ld=c. If ld=c and aj+bk=d, then aj+bk=c/l \[aj+bk=\frac{c}{l}=>ajl+bkl=c\] but ji and kl are integers so let's just write M for ji and L for kl where M and L are integers. So we have \[aM+bL=c\] by equation (1) this implies \[M=m^2 \& L=n^2\] if M and L are not perfect squares, then we do not have integers if M and L are perfect squares, then we do have integers m and n

OpenStudy (anonymous):

but if m is an integer then so is m^2

myininaya (myininaya):

but if M is not a perfect square and if L is not a perfect square then m and n are not integers (contradiction)

myininaya (myininaya):

and lets say or instead of and

OpenStudy (anonymous):

ok so you have disproven it thanks

myininaya (myininaya):

but if M is not a perfect square or if L is not a perfect square then m or n are not integers (contradiction)

myininaya (myininaya):

we it is true for perfect squares M and L

myininaya (myininaya):

so yes i have disproved it

myininaya (myininaya):

since it is not true for all M and L

OpenStudy (anonymous):

o right because its "there exists" not "for all"

OpenStudy (anonymous):

man im confused

OpenStudy (anonymous):

i think now youve proven it

myininaya (myininaya):

lol i think to because i did prove there exists integers that satisfy that equation given the conditions

OpenStudy (anonymous):

yep ok ill work on it thanks

OpenStudy (zarkon):

for all a,b,c are integers, there exists an integer solution to ax^2 +by^2=c, whenever gcd(a,b)|c for any integers ...I pick a=2,b=3 and c=1 then the gcd(a,b)=1 and 1|c but 2x^2 +3y^2=1 has no integer solutions

myininaya (myininaya):

oh yes it is easier than i thought just find a counterexample

myininaya (myininaya):

zarkon what do you think of what i wrote?

OpenStudy (anonymous):

here is what i got

myininaya (myininaya):

whats my grade?

OpenStudy (zarkon):

jas...I'm switching browsers

OpenStudy (anonymous):

d=gcd(a,b) d|c kd=c ax+by=d so ax+by=c/k then akx +bky=c this implies that kx=x^2, and ky=y^2 this can't be true because y cant equal x

OpenStudy (zarkon):

I would just stick to a counter example

OpenStudy (anonymous):

ya but they say prove i think they need that

OpenStudy (zarkon):

@myininaya looks good...I can't give an A+ though since you did it the long way :) Also it is possible the M and L are perfect squares (some of the numbers do have solutions)

myininaya (myininaya):

:(

myininaya (myininaya):

right

myininaya (myininaya):

and i stated this so i get an A

myininaya (myininaya):

A+*

OpenStudy (zarkon):

the Dual of A+...sweet!

OpenStudy (anonymous):

i got a lot of duals of A+ i think i lived in the dual space all through school

myininaya (myininaya):

lol

OpenStudy (zarkon):

sorry I couldn't resist http://en.wikipedia.org/wiki/Dual_space

myininaya (myininaya):

satellite someone thought we were bad earlier

OpenStudy (anonymous):

really? i thought i deleted everything bad that i wrote

OpenStudy (anonymous):

if gcd(a,b)=1, then gcd (2a+b,a+2b) is 1 or 3 : prove

OpenStudy (anonymous):

@myininaya can you tell me in chat?

OpenStudy (zarkon):

ok....notice that \[2(2a + b) − (a + 2b) = 3a.\] \[2(a + 2b) − (2a + b) = 3b\] now work it out from here ...

OpenStudy (anonymous):

k i know where to go from there but did you just get the 2 intuitively?

OpenStudy (anonymous):

and why minus

OpenStudy (zarkon):

I was just playing around with linear combinations...I should have seen it much sooner than I did :)

OpenStudy (anonymous):

does d|any linear combination of 2a+b and 2b+a

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

then d=gcd(3a,3b) d=3gcd(a,b) then 3d=gcd(a,b)

OpenStudy (zarkon):

not exactly

OpenStudy (anonymous):

d/3

OpenStudy (zarkon):

d|gcd(3a,3b)=3gcd(a,b)=3 thus d|3

OpenStudy (anonymous):

right and the 1 is given

OpenStudy (zarkon):

yep

OpenStudy (anonymous):

k cool thx